OpenStudy (anonymous):

Find the geometric mean of square root of 28 and square root of 7.

OpenStudy (ash2326):

Geometric mean of two numbers a and b is given as \[\large GM=\sqrt{ab}\] here \(a= \sqrt {28}\) and \(b= \sqrt 7\) so \[GM=\sqrt {\sqrt{28} \times \sqrt {7}}\] We have now \[\large GM=\sqrt {\sqrt{196}}\] or \[\large GM =\sqrt{14}\] The GM is \[\huge \sqrt {14}\]